OpenStudy (anonymous):

Is my answer right??? please check! thanks! 6x^2 - 54 + 84 over 8x^2 - 40x + 48 + x^2 + x - 56 over 2x^2 + 12x - 32 My answer: 2(x - 2) over 3(x - 3)

OpenStudy (anonymous):

@jim_thompson5910

sam (.sam.):

Hmm, is it 6x^2 - 54 + 84 over 8x^2 - 40x + 48 = x^2 + x - 56 over 2x^2 + 12x - 32 ?

jimthompson5910 (jim_thompson5910):

The problem is \[\Large \frac{6x^2 - 54 + 84}{8x^2 - 40x + 48} + \frac{x^2 + x - 56}{2x^2 + 12x - 32}\] right?

OpenStudy (anonymous):

yes sir

jimthompson5910 (jim_thompson5910):

what you got is incorrect

jimthompson5910 (jim_thompson5910):

How did you get that?

OpenStudy (anonymous):

Your answer is right but you must rotate 3(x - 2)/(2(x - 3))

OpenStudy (anonymous):

thank you so so so much sir

OpenStudy (anonymous):

YW !